Please help me solve this, thanks!

Find the general solution to the system x = Ax where A is the given matrix. | -2 -2 -6 A= 0 0 6 | 0 -2 -8 b) X(t)=( X(t)= Ce0 e) X(t)= C, e 20 +46? -6 +2° -1 | 2 f) None of the above.Find the general solution to the system x= Ax where A is the given matrix. 0 1 0 A= 0 0 1 | -20 16 -1 +te b) XO = ( 27 -51 cf) None of the above.
Answer #1

$ 01 AUGUST : SATURDAY AUGUST. SATURDAY 19 20 222 75 1) Here X = AX net A= 1-2-2-6 Lor2-8 Post of we that find eigen valuesMONDAY. AUGUST MONDAY • AUGUST 03 OC o -2-6 22 O-2- C1710 - fah ZR & R Pz - BRT R2 R₂ + R & R ₂ - R &R o-2-6 Toy 22 Looola loAUGUST TUESDAY Ola eigen vector corresponding to eigen value put d =-2 9=2 8-2 care Tol 2 To vector corresponding to find eigS M T WT 31 JULYO FRIDAY JULY FRIDAY 29 30 21-1930 4- 2 - 6T 06 6 Oo o 23 system has infinite solh take ( =/ x2 = -1 4, -2726FRIDAY. AUGUST WK 32) 220-146 1-2016-1 Coulder | Adl=0 - 10 O- 1 1-20 16 -1-d - [(2+) -16] - (20) -[d?td-16 ]-20 d3d²+ 16d-20S M T W T F 08 S AUGUST SATURDAY R3 R3TIORI P R + 8 R2 system has infinite son put x =P 2,= ply SUNDAY 22 2020 2020-4-27 00:1MONDAY O RUCUST 10 MONDAY • AUGUST eigen vector for di-5 | A-da [x]=ro] 1-20 16 4 02 223 R2 Z R3t4RI az To 204 Rqg R3-4R2 0 5M T W T F S AUGUST TUESDAY ME un roots are repeated to Xl 5H col T-5 I 25 2020 2020-4-27 00:10